OpenStudy (buggiethebug):

determine the derivative: y=(3-2x^3)^3 can someone show me the steps? I dont know what I'm doing wrong

OpenStudy (anonymous):

Make sure you're using the chain rule: 1) bring down the exponent in front and subtract 1 from it\[y'=3(3-2x^3)^2\] but don't forget to multiply that by the derivative of the inside of the parenthesis. The derivative of :\[3-2x^3\] is \[-6x^2\] so multiply the original derivative we got \[y'=3(3-2x^3)^2\] by the derivative of the inside of the parenthesis\[-6x^2\] to get \[y'=3(3-2x^3)^2(-6x^2)=-18 x^2 (3-2 x^3)^2\]

OpenStudy (buggiethebug):

but we're only supposed to use the product rule, sorry that i forgot to mention

OpenStudy (anonymous):

\[y=(3-2x^3)^3\] is the same as \[y=(3-2x^3)(3-2x^3)^2\] Square out the one that's squared: \[(3-2x^3)^2=4 x^6-12 x^3+9\] Giving you \[y=(3-2x^3)(4 x^6-12 x^3+9)\] Product rule gives you \[y'=(3-2x^3)(24 x^5-36 x^2)+(4 x^6-12 x^3+9)(-6x^2)\] a little simplified: \[y'=(-48 x^8+144 x^5-108 x^2)+(-24 x^8+72 x^5-54 x^2)\] and more: \[y'=-48 x^8+144 x^5-108 x^2-24 x^8+72 x^5-54 x^2)\] and finally:\[y'=-72 x^8+216 x^5-162 x^2\] simplifying to:\[y'=-18 x^2 (2 x^3-3)^2\] Which is the same as the answer I originally posted:\[y'=-18 x^2 (3-2 x^3)^2\] .

OpenStudy (precal):

you will later find out that chain rule is faster then product rule in this specific case

OpenStudy (buggiethebug):

thank you!